OpenStudy (anonymous):

Differential equation. (x+y)y'=1

OpenStudy (anonymous):

Multiply both sides by dx, and what do you get?

OpenStudy (anonymous):

(x+y)dy=dx

OpenStudy (anonymous):

Integrate both sides.

OpenStudy (anonymous):

xy+y^(2)/2=x+c

OpenStudy (anonymous):

forgot the dy

OpenStudy (anonymous):

This is now a simple problem.

OpenStudy (anonymous):

(x+y) = y'

OpenStudy (anonymous):

Woo.